OpenStudy (anonymous):

write a rule for the linear function in the table. (x) -4,0,4,8 / f(x) 16,0,-16,-32

OpenStudy (anonymous):

answers : A) f(x)=-4x B) f(x)=x+16 C) f(x)=4x D) f(x)=-1/4x